Novell, Inc., has the following mutually exclusive projects.

  

Year Project A Project B
  0 –$21,000    –$24,000   
  1 12,500    13,500   
  2 9,000    10,000   
  3 3,000    9,000   

  

a-1.

Calculate the payback period for each project. (Do not round intermediate calculations and round your answers to 3 decimal places, e.g., 32.161.)

a-2.

If the company's payback period is two years, which, if either, of these projects should be chosen?

  • Project A

  • Project B

  • Both projects

  • Neither project

  

b-1.

What is the NPV for each project if the appropriate discount rate is 17 percent? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

b-2.

Which, if either, of these projects should be chosen if the appropriate discount rate is 17 percent?

  • Project A

  • Project B

  • Both projects

  • Neither project
Answer #1

For calculating payback period we are using excel table

Year 0 1 2 3
Project A cash flows -21000 12500 9000 3000
Cumulative cash flow -21000 -8500 500 3500
Payback Period 1.94 Payback Period = 1+8500/9000
Year 0 1 2 3
Project B cash flows -24000 13500 10000 9000
Cumulative cash flow -24000 -10500 -500 8500
Payback Period 2.06 Payback Period = 2+500/9000

a-1) Payback period of A = 1.94
Payback period of B = 2.06

a-2) Project A should be selected because payback period is less than 2 years.

b-1) NPV of Project A = - 21000 + 12500/(1+17%) + 9000/(1+17%)2 + 3000/(1+17%)3 = -1868.51
NPV of Project B = - 24000 + 13500/(1+17%) + 10000/(1+17%)2 + 9000/(1+17%)3 = 462.93

b-2) Project B should be selected.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well